OpenStudy (anonymous):

Use the quadratic formula to solve x^2 + 7x + 8 = 0. Estimate irrational numbers to the nearest tenth. A. {-12, 5} B.{-11.1, -2.9} C. {-5.6, -1.4} D. {-3.9, 12.1}

OpenStudy (photon336):

We need to use this formula here \[-b \pm \frac{ \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (photon336):

\[x^{2}+7x+8 = 0\] \[ax^{2}+bx+c\] so basically a would be the number in front of x^{2} b is the number in front of x. and c is the last number so a would be 1 b would be 7 and c 8 try plugging these numbers into the equation and then tell me what you get